The fundamental and 1st overtone of a molecule are found at 2886 cm^-1 and 5668 cm^-1 respectively. Calculate the equilibrium oscillation wave no. and the anharmonicity constant.
1
Expert's answer
2019-05-29T05:16:41-0400

V1-V2=2886 - 5668 = 2782.

1/2782